# Non-Asymptotic Analysis of Efficiency in Conformalized Regression

### v1

@@ Appendix E Experiments on Real-World Data
The Medical Expenditure Panel Survey (MEPS) Panels 19111https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-181 and 20222https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-192 are standard datasets used for benchmarking and comparative analysis in the quantile regression literature. These panels comprise 15,785, 17,541, and 15,656 samples, respectively. Each sample consists of 139 features, including 2 categorical features, 4 continuous features, and 133 boolean features. Throughout experiments, we train ridge regression models with stochastic gradient descent (SGD) optimizer with a step size tuned using successive halving for 1 epoch.
Conformalized Median Regression (CQR).
@@
We examine the effect of the training set size $n$ and the calibration set size $m$ on the prediction set length using the MEPS’20 dataset, comparing the empirical results with the theoretical bound in Theorem 3.2. Since the oracle quantile interval length $|\mathcal{C}^{*}(X)|=q_{1-\alpha/2}(Y|X)-q_{\alpha/2}(Y|X)$ depends on $\alpha$, we evaluate the expected absolute deviation $\mathbb{E}[||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)||]$ for $\alpha\in[0.01,0.05,0.1,0.2]$. We reserve 20% of the dataset for testing length deviation. The remaining 80% was partitioned for training and calibration: the training size $n$ varied from 10% to 80% in increments of 10%, while the calibration $m$ was chosen from $5\%,10\%,15\%,20\%$} of the remaining data after allocating the training set. The results, shown in Fig. 6, confirm two key insights from our theoretical analysis. First, increasing the calibration set size $m$ reduces the expected length deviation. Second, for a fixed sample size, a larger miscoverage level $\alpha$ leads to a smaller deviation with lower variance, which aligns with the $\alpha$-dependence in the theoretical rate.
Conformalized Median Regression (CMR).
Figure 7 presents experimental results on length deviation for the MEPS’19 dataset under the CMR framework. The experimental setup mirrors that of the previous CQR analysis, adapted here for median regression. Consistent with Theorem 4.1, we observe that smaller values of $\alpha$ yield significantly larger length deviations.
Figure 6: The efficiency of CQR. Here the y-axis represents $\left||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)|\right|$, where the interval length $|\mathcal{C}^{*}(X)|$ is approximated by its estimate with same $\alpha$ and largest training and calibration sample sizes.
Figure 7: The efficiency of CMR. Here the y-axis represents $\left||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)|\right|$, where the interval length $|\mathcal{C}^{*}(X)|$ is approximated by its estimate with same $\alpha$ and largest training and calibration sample sizes.

### v2

@@ D.1 Statistics of Datasets
The Medical Expenditure Panel Survey (MEPS) Panels 19111https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-181 and 20222https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-192 are standard datasets used for benchmarking and comparative analysis in the quantile regression literature. Each sample consists of 139 features, including 2 categorical features, 4 continuous features, and 133 boolean features.
Table 1: Statistics of datasets.
Dataset # Features # Number Samples
MEPS 19 139 15,785
MEPS 20 139 17,541
cpusmall (CC01a) 12 8,192
abalone (CC01a) 8 4,177
California Housing (pace1997sparse) 8 20,640
D.2 Empirical Evaluation of Length Deviation
D.2.1 Experimental Settings
@@
We examine the effect of the training set size $n$ and the calibration set size $m$ on the prediction set length, comparing the empirical results with the theoretical bound in Theorem 3.2. Since the oracle quantile interval length $|\mathcal{C}^{*}(X)|=q_{1-\alpha/2}(Y|X)-q_{\alpha/2}(Y|X)$ depends on $\alpha$, we evaluate the expected absolute deviation $\mathbb{E}[||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)||]$ for $\alpha\in[0.01,0.05,0.1,0.2]$, where the interval length $|\mathcal{C}^{*}(X)|$ is approximated by its estimate with same $\alpha$ and largest training and calibration sample sizes. We reserve 20% of the dataset for testing length deviation. The remaining 80% data was partitioned for 80% training data and 20% calibration data: the training size $n$ varied from 10% to 80% in increments of 10%, while the calibration $m$ was chosen from $5\%,10\%,15\%,20\%$} of the remaining data. Throughout experiments, models are trained with a step size tuned by successive halving for 1 epoch.
D.2.2 Empirical Results with Various Optimizers
Figure 11 presents an empirical evaluation of length deviation of CMR and CQR under different optimizers on real-world datasets, comparing SGD, SGD with momentum (SGDM) (polyak1964some), Adam (kingma2015adam), and AdamW (loshchilov2019decoupled). Although our theory is based on analyzing SGD and directly extends to SGD with momentum, we include Adam and AdamW due to their widespread practical use.
The results confirm two key insights from our theoretical analysis. First, increasing the calibration set size $m$ reduces the expected length deviation. Second, for a fixed sample size, a larger miscoverage level $\alpha$ leads to a smaller deviation with lower variance, which aligns with the $\alpha$-dependence in the theoretical rate. Consistent with Theorem 4.1, we observe that smaller values of $\alpha$ yield significantly larger length deviations.
Among the optimizers, we observe that Adam and AdamW generally achieve better efficiency (lower length deviation) but exhibit higher volatility, likely due to their scaled gradient norms. SGD decays more smoothly, providing a more consistent reduction in length deviation as the number of training samples increases. For Adam and AdamW, the benefit of additional training data can be less pronounced or even reversed on certain datasets (e.g., MEPS), where fast convergence makes efficiency more sensitive to stochasticity. On other datasets such as abalone and cpusmall, however, Adam also exhibits a clear decreasing trend, indicating a dataset-dependent behavior.
(a) CMR
(b) CQR
Figure 11: Efficiency of conformalized regression under different training optimizers on real-world datasets: MEPS 19, MEPS 20, California Housing (pace1997sparse), cpusmall, and abalone (CC01a). For each optimizer, the learning rate is selected via successive halving, while all other hyperparameters (e.g., momentum=0.9 for SGD with momentum) follow the PyTorch defaults.
D.2.3 Empirical Probing on Non-Linear Models
We conduct empirical probing of non-linear models on real-world datasets, and report the results in Figure 12. We observe that the length deviation remains consistent across non-linear and linear model architectures, suggesting a potential practical relevance of our findings beyond linear models.
(a) CMR
(b) CQR
Figure 12: Efficiency of conformalized regression with linear and non-linear models trained via SGD on real-world datasets: MEPS 19, MEPS 20, California Housing (pace1997sparse), cpusmall, and abalone (CC01a). The two-layer neural network has one hidden layer with 10 ReLU neurons, and the three-layer network has two hidden layers, each with 10 ReLU neurons.
D.3 Empirical Data Allocation Guidance
We empirically investigate how to allocate data on cpusmall dataset from LIBSVM (CC01a). The training ratio $r_{n}$ takes values from [0.01, 0.05, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 0.95, 0.99], the calibration ratio $r_{m}$ is set as $1-r_{n}$, the miscoverage level $\alpha$ takes values from [0.001, 0.002, 0.003, 0.004, 0.005, 0.006, 0.007, 0.008, 0.009, 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.15, 0.2].
The left plot in Figure 13 shows the length of the prediction interval versus $\alpha$ of CMR, grouping curves by training ratio (0%–20%, 20%–80%, 80%–100%). We observe that two “elbows” occur at approximately $\alpha=0.045$ and $\alpha=0.003$, at which points, reducing $\alpha$ leads to a substantially sharper increase in interval length than before. Notably, before the first elbow, e.g., when reducing $\alpha$ from $0.2$ to $0.05$, the prediction interval length increases only mildly.
The right plot in Figure 13 shows the length of the prediction interval versus the training ratio, with each curve corresponding to a different miscoverage level $\alpha$ (lighter color representing smaller $\alpha$). We observe that:
  • The curves largely concentrate around interval lengths of approximately $2.5$ and $15$, respectively, which correspond to the two elbow locations in the left plot.
  • For most cases where $\alpha$ is not extremely small, the interval length stays below $15$, and the curves exhibit a wide U-shape. This indicates that allocating an excessively large portion of data to either training or calibration tends to degrade efficiency, whereas a more balanced split yields better efficiency. For reasonably large $\alpha$, say $\alpha>0.04$, the number of calibration samples has less influence on the interval than the number of training samples, suggesting that allocating more data for training is generally beneficial.
  • For very small miscoverage levels ($\alpha\leq 0.003$), corresponding to the three curves above the dashed line at length =15=15, the interval length behaves erratically and no longer follows the U-shaped trend observed for larger $\alpha$. This is likely due to insufficient sample size at such small $\alpha$, and the prediction interval length becomes a trivial upper bound of the oracle interval length rather than its approximation. This phase of extremely small $\alpha$ may correspond to the regime $\alpha=\omega(n^{-1/2})$, where our upper bound is non-vanishing (Figure 2).
Takeaway. The empirical results suggest that practitioners may leverage the elbow points in the left plot of Figure 13 to select $\alpha$ values that yield good efficiency while maintaining reasonable coverage guarantees. In particular, for extremely small $\alpha$, the prediction interval becomes trivially large due to insufficient sample size, whereas in the regime of large $\alpha$, decreasing the miscoverage level results in only a mild increase in interval length. In terms of data allocation, the results are consistent with the practical rule of thumb that the amount of training and calibration data should be roughly of the same order, while allocating slightly more data for training is generally beneficial.
Figure 13: The effect of training ratio $r_{n}$, calibration ratio $r_{m}$, and miscoverage level $\alpha$ on cpusmall dataset. The training ratio $r_{n}$ takes values from $0.01$ to $0.99$, the calibration ratio $r_{m}$ is $1-r_{n}$, and $\alpha$ takes values from $0.001$ to $0.2$. See Appendix D.3 for detailed discussion on empirical data allocation. The left plot shows the length of the prediction interval versus $\alpha$, grouping curves by training ratio (0%–20%, 20%–80%, 80%–100%). We observe that there are two “elbows” around $\alpha=0.045$ and $\alpha=0.003$, at which points, reducing $\alpha$ leads to sharper rise of the interval length than before. The right plot shows the length of the prediction interval versus the training ratio, with each curve corresponding to a different miscoverage level $\alpha$ (lighter color representing smaller $\alpha$).
D.4 Distribution Patterns of Real-World Datasets
To illustrate the distribution patterns of the real-world datasets used in our experiments, Figure 14 visualizes the distributions of the first 90 features of MEPS 19, and Figure 15 summarizes all feature distributions of the California Housing dataset. For completeness, the abalone and cpusmall datasets are also shown in Figures 17 and 18.
To investigate heavy-tailed features, we use the California Housing dataset as a case study. We apply the Hill estimator (hill1975simple) to compute the extreme value index $\xi$ and systematically detect heavy-tailed features. Following the framework of voitalov2019scale, we conservatively classify features with $\xi>1/4$ as heavy-tailed, identifying three features total_rooms, total_bedrooms, and population. This classification is supported by the numerical estimates in Figure 15 and the QQ plots in Figure 16.
With this investigation of the heavy-tailed features, the experimental results in Figure 11 and Figure 12 show that the length deviation in general behaves consistently among datasets, suggesting that CQR and CMR remain robust even in the presence of heavy-tailed features.
Figure 14: Distribution patterns of features in the MEPS 19 dataset, displaying the initial 90 features from the complete set of 139 variables
Figure 15: Histograms of features from the California Housing Dataset (pace1997sparse). The tail index $\xi$, estimated via the Hill estimator (hill1975simple), characterizes the maximum domain of attraction (MDA). A value of $\xi=0$ corresponds to the Gumbel MDA (light-tailed), while $\xi>0$ indicates the Fréchet MDA (heavy-tailed). Following the conservative threshold suggested by voitalov2019scale ($\xi>0.25$), we identify three features as heavy-tailed: total rooms, total bedrooms, and population. The features have been standardized.
Figure 16: $Q$-plot of the features in the California Housing Dataset (pace1997sparse).
Figure 17: Histograms of features from the abalone dataset (CC01a). Settings are the same as Figure 15.
Figure 18: Histograms of features from the cpusmall dataset (CC01a). Settings are the same as Figure 15.